A jet lands on a aircraft carrier at 84 mile (33m/s).
(a)what is its acceleration if it stops in 27 sec.
(b)what is the displacement of the plane which it is stopping.
1
Expert's answer
2020-03-04T10:02:35-0500

Given:

"v_i=33\:\rm m/s;"

"v_f=0\:\rm m/s;"

"t=27\:\rm s."

(a) The acceleration


"a=\frac{v_f-v_i}{t}=\frac{0-33}{27}=-1.22\:\rm m/s^2."

(b) The displacement


"d=\frac{v_f+v_i}{2}t=\frac{0+33}{2}\times 27=445.5\:\rm m."